On the distant planet Cowabunga , the weights of cows have a normal distribution with a mean of 357 pounds and a standard deviation of 48 pounds. The cow transport truck holds 10 cows and can hold a maximum weight of 3910. If 10 cows are randomly selected from the very large herd to go on the truck, what is the probability their total weight will be over the maximum allowed of 3910? (This is the same as asking what is the probability that their mean weight is over 391.)

(Give answer correct to at least three decimal places.)
probability = On the distant planet Cowabunga , the weights of cows have a normal distribution with a mean of 357 pounds and a standard dev
Answer #1

Solution

\mathbf{given:}\;\mu=357,\;\sigma=48,\;n=10

\mathbf{formula:}\;z=\frac{\bar x-\mu}{\frac{\sigma }{\sqrt{n}}}

P(mean\;weight\;is\;over\;391)\Rightarrow P(\bar x>391)

\Rightarrow P\left (\frac{\bar x-\mu}{\frac{\sigma }{\sqrt{n}}}>\frac{391-357}{\frac{48}{\sqrt{10}}} \right )

\Rightarrow P\left (z>2.24 \right )

\mathbf{Note}:\;P(z>a)=1-P\left (z<a\right )

\Rightarrow 1-P\left (z<2.24 \right )

Refer Z-table to find the probability or use excel formula "=NORM.S.DIST(2.24, TRUE)" to find the probability.

\Rightarrow 1-0.9875

\Rightarrow \mathbf{{\color{Red} 0.0125}}

\mathbf{\therefore probability= \mathbf{{\color{Red} 0.0125}}}